# BiasBusters: Uncovering and Mitigating Tool Selection Bias in Large Language Models

### v1

@@ Appendix E More Elaborate Ablation and Sensitivity Analysis
Temperature. Raising temperature reduces combined bias. As shown in Figure 9(a), as temperature goes from 0 to 2, the mean $\delta_{\mathrm{model}}$ for ChatGPT 3.5 drops from about 0.350 to 0.285, a 6.5% absolute reduction. Figure 10 makes clear why: the overall selection patterns remain similar across temperatures, but higher temperatures soften extreme preferences. This suggests that increased stochasticity slightly mitigates bias, but does not eliminate it.
Top-p. Figure 9(b) shows how the combined bias $\delta_{\mathrm{model}}$ for ChatGPT 3.5 varies with the top-p cutoff. Increasing top-p from 0.7 to 1.0 yields a small decrease in bias (from $\sim$0.346 to $\sim$0.340), suggesting that less aggressive truncation of the probability distribution slightly softens extreme tool preferences. The effect is noticeably weaker than the temperature change.
@@
API Ordering. Figure 11 compares ChatGPT 3.5’s API selection under two different ordering schemes: cyclic rotations versus random permutations. Across all clusters, the choice distribution is very similar: no API’s selection rate shifts more than about ten percentage points. This indicates that the ordering of the APIs has some influence, but the dominant signal is the model’s intrinsic preference. However, the small differences could also reflect the inherent noise from stochastic token sampling, and overall we argue that the tool-selection behavior is robust to either type of shuffling.
System Prompts. To evaluate how sensitive tool selection is to the phrasing and structure of the instructions given, we compare three variants of the system prompt: the original “Base” prompt, a lightly reworded “Similar” prompt, and a structurally different “Adjusted” prompt. Figure 12 shows the resulting distributions for ChatGPT 3.5.
@@
Figure 10: Selection distributions for ChatGPT 3.5 using four different temperatures across ten clusters of functionally equivalent APIs. Each subplot corresponds to one cluster, with the x-axis indicating the API in the cluster and the y-axis showing the fraction of times that API was chosen by the respective model over 500 runs.
Figure 11: Selection distributions for ChatGPT 3.5 using cyclic and random shuffling of the APIs across ten representative clusters. Each subplot corresponds to one cluster, with the x-axis indicating the API in the cluster and the y-axis showing the fraction of times that API was chosen for that specific API ordering method over 500 runs.
Figure 12: Selection distributions for ChatGPT 3.5 using different system prompts across ten clusters. Each subplot corresponds to one cluster, with the x-axis indicating the API in the cluster and the y-axis showing the fraction of times that API was chosen using the corresponding system prompt over 500 runs.
@@ F.1 Which API-level features predict selection rates?
Similarity between tool / API description and query is most correlated to selection rate. As Table 4 makes clear, the most predictive feature of API selection is semantic similarity between the query and the tool/API descriptions. Both avg_similarity_tool_desc and avg_similarity_api_desc are consistently positively correlated with selection rates—especially strong for Qwen and clear for Gemini; ChatGPT shows the same pattern, albeit weaker with higher p-values. By contrast, structural or stylistic attributes (e.g., parameter count, promotional wording) exhibit little consistent influence. Tool age (age_days) shows a modest, broadly consistent negative correlation across models.
@@
Feature ChatGPT 4.1 Gemini Qwen
@@
Linear regression reveals that surface-level semantic alignment is the primary signal but leaves a lot unexplained. Figure 13 tells us that linear models explain only part of the variance: $R^{2}$ is modest—0.143 for ChatGPT-4.1 and 0.387 for ToolLLaMA—leaving substantial error. Coefficients show surface-level semantic alignment dominates: similarity between the query and tool/API descriptions has the largest positive weights for most models, with Qwen weighting both most strongly and Gemini emphasizing API-level descriptions. Unexpectedly, ToolLLaMA gives a negative weight to tool-description similarity. Other features contribute little. Hence, semantic alignment is important in driving selection but still gives an incomplete explanation, implying nonlinear or omitted factors and motivating more flexible models (e.g., random forests).
@@ Appendix H Effect of Metadata Perturbation on Bias
Relative to the base distributions, both models move farther from uniform (get more biased) when we lower semantic signal (see bars corresponding to the description + parameter and full perturbations in Figure 18). For Gemini, these manipulations yield the largest TV distances to uniform ($\approx$0.42–0.43); ChatGPT shows similar results. This suggests that when descriptions/parameters are corrupted, models amplify bias rather than flatten choices.
@@
Figure 18: Mean total-variation (TV) distance from the uniform selection distribution to the distribution pertaining to each metadata perturbation (higher = more deviation from uniform). Blue bars show Gemini and orange bars show ChatGPT; error bars denote standard deviation across clusters, and single bars indicate perturbations not run for ChatGPT. Dashed horizontal lines (in the corresponding model colors) mark each model’s baseline TV-to-uniform without perturbations.

### v2

@@ Appendix E More Elaborate Ablation and Sensitivity Analysis
Temperature. Raising temperature reduces combined bias. As shown in Figure 9(a), as temperature goes from 0 to 2, the mean $\delta_{\mathrm{model}}$ for GPT-3.5 drops from about 0.350 to 0.285, a 6.5% absolute reduction. Figure 10 makes clear why: the overall selection patterns remain similar across temperatures, but higher temperatures soften extreme preferences. This suggests that increased stochasticity slightly mitigates bias, but does not eliminate it.
Top-p. Figure 9(b) shows how the combined bias $\delta_{\mathrm{model}}$ for GPT-3.5 varies with the top-p cutoff. Increasing top-p from 0.7 to 1.0 yields a small decrease in bias (from $\sim$0.346 to $\sim$0.340), suggesting that less aggressive truncation of the probability distribution slightly softens extreme tool preferences. The effect is noticeably weaker than the temperature change.
@@
API Ordering. Figure 11 compares GPT-3.5’s API selection under two different ordering schemes: cyclic rotations versus random permutations. Across all clusters, the choice distribution is very similar: no API’s selection rate shifts more than about ten percentage points. This indicates that the ordering of the APIs has some influence, but the dominant signal is the model’s intrinsic preference. However, the small differences could also reflect the inherent noise from stochastic token sampling, and overall we argue that the tool-selection behavior is robust to either type of shuffling.
System Prompts. To evaluate how sensitive tool selection is to the phrasing and structure of the instructions given, we compare three variants of the system prompt: the original “Base” prompt, a lightly reworded “Similar” prompt, and a structurally different “Adjusted” prompt. Figure 12 shows the resulting distributions for GPT-3.5.
@@
Figure 10: Selection distributions for GPT-3.5 using four different temperatures across ten clusters of functionally equivalent APIs. Each subplot corresponds to one cluster, with the x-axis indicating the API in the cluster and the y-axis showing the fraction of times that API was chosen by the respective model over 500 runs.
Figure 11: Selection distributions for GPT-3.5 using cyclic and random shuffling of the APIs across ten representative clusters. Each subplot corresponds to one cluster, with the x-axis indicating the API in the cluster and the y-axis showing the fraction of times that API was chosen for that specific API ordering method over 500 runs.
Figure 12: Selection distributions for GPT-3.5 using different system prompts across ten clusters. Each subplot corresponds to one cluster, with the x-axis indicating the API in the cluster and the y-axis showing the fraction of times that API was chosen using the corresponding system prompt over 500 runs.
@@ F.1 Which API-level features predict selection rates?
Similarity between tool / API description and query is most correlated to selection rate. As Table 4 makes clear, the most predictive feature of API selection is semantic similarity between the query and the tool/API descriptions. Both avg_similarity_tool_desc and avg_similarity_api_desc are consistently positively correlated with selection rates—especially strong for Qwen and clear for Gemini; GPT shows the same pattern, albeit weaker with higher p-values. By contrast, structural or stylistic attributes (e.g., parameter count, promotional wording) exhibit little consistent influence. Tool age (age_days) shows a modest, broadly consistent negative correlation across models.
@@
Feature GPT-4.1 mini Gemini Qwen
@@
Linear regression reveals that surface-level semantic alignment is the primary signal but leaves a lot unexplained. Figure 13 tells us that linear models explain only part of the variance: $R^{2}$ is modest—0.143 for GPT-4.1 mini and 0.387 for ToolLLaMA—leaving substantial error. Coefficients show surface-level semantic alignment dominates: similarity between the query and tool/API descriptions has the largest positive weights for most models, with Qwen weighting both most strongly and Gemini emphasizing API-level descriptions. Unexpectedly, ToolLLaMA gives a negative weight to tool-description similarity. Other features contribute little. Hence, semantic alignment is important in driving selection but still gives an incomplete explanation, implying nonlinear or omitted factors and motivating more flexible models (e.g., random forests).
@@ Appendix H Effect of Metadata Perturbation on Bias
Relative to the base distributions, both models move farther from uniform (get more biased) when we lower semantic signal (see bars corresponding to the description + parameter and full perturbations in Figure 18). For Gemini, these manipulations yield the largest TV distances to uniform ($\approx$0.42–0.43); GPT shows similar results. This suggests that when descriptions/parameters are corrupted, models amplify bias rather than flatten choices.
@@
Figure 18: Mean total-variation (TV) distance from the uniform selection distribution to the distribution pertaining to each metadata perturbation (higher = more deviation from uniform). Blue bars show Gemini and orange bars show GPT; error bars denote standard deviation across clusters, and single bars indicate perturbations not run for GPT. Dashed horizontal lines (in the corresponding model colors) mark each model’s baseline TV-to-uniform without perturbations.
@@ Appendix I Details on CPT Setup
Appendix J Effect of Tool Count on Bias
To investigate how the number of available tools affects selection bias, we conducted an additional experiment using the Sentiment Analysis cluster (which contains five functionally equivalent APIs). From this cluster, we constructed subsets of size $K\in\{2,3,4\}$ using three selection strategies:
  • Best-to-worst: select the $K$ APIs that were most frequently chosen in our initial $K=5$ experiments.
  • Worst-to-best: select the $K$ least frequently chosen APIs.
  • Random subsets: uniformly sample $K$ tools from the cluster.
For each subset configuration, we re-ran the corresponding queries three times using Qwen3 (235B) and computed the normalized API-level and positional bias metrics.
Figure 19: Normalized API-level and positional bias as a function of cluster size $K$ for the sentiment-analysis cluster. Each curve corresponds to one subset-selection strategy (best-to-worst, worst-to-best, random), with error bars indicating variability across three independent runs. Bias is reported as a percentage of its theoretical maximum (11/K1-1/K). The rightmost subplot shows the average of the normalized API-level and positional bias over $K$.
As can be seen in Figure 19, we find that bias is highly sensitive to the specific subset of tools shown to the model. At $K=2$, both normalized $\delta_{\text{API}}$ and $\delta_{\text{pos}}$ vary substantially across selection strategies, revealing strong instability. At moderate subset sizes ($K=3,4$), the variance decreases and the measured bias becomes more stable, although not uniformly smaller. Overall, the relationship between $K$ and bias is non-monotonic: depending on which tools are included, reducing the number of available tools can either amplify or attenuate bias. These findings indicate that the composition of the toolset is the primary driver of selection behavior, more than its size alone.
Finally, note that when $K=5$ all methods trivially converge, since they expose the full cluster and therefore share identical toolsets. To fully characterize tool-count effects, future work should extend this analysis to larger clusters (e.g., K=6K=61010).
Appendix K Explanation of LLM Tool Usage Pipeline
Figure 20: The overall workflow for tool usage with large language models. This figure illustrates the last three stages of tool usage: tool selection (via retriever and LLM selection), tool calling, and response generation (task planning is omitted). Note how the LLM’s preliminary outputs need to be parsed to create a functional API call.
Large language models equipped with external tools typically follow a multi-stage decision and execution process. As summarized by prior work (Qu et al., 2025), the standard tool-usage pipeline consists of four stages: task planning, tool selection, tool calling, and response generation. Figure 20 showcases the latter three of these stages.
In the first stage, the LLM interprets the user query, identifies the user’s intent, and (if necessary) decomposes it into sub-tasks that can be handled by external tools. Next, in the tool selection stage, the model determines which tools are suitable for each sub-task. Modern systems often employ a retriever-based filtering step before LLM-based selection: rather than presenting the descriptions of hundreds or thousands of available tools to the model (which is the case with large API hubs like RapidAPI), a retriever identifies the top-k most relevant candidates. This design is widely adopted in practical deployments due to context-length limitations (Qu et al., 2025).
In the tool calling stage, the LLM chooses one of the retrieved tools and generates the corresponding API call, including the required parameters in the correct format. This step requires the model not only to select an appropriate tool but also to extract and structure the tool arguments accurately. Once the LLM generation is parsed, the tool executes and returns its output, the LLM may decide to proceed to the final stage or determine that more tools need to be called. In the final stage, the LLM integrates the returned tool results into a coherent final answer to the user.
Appendix L Consequences of Tool Selection Bias
We have briefly touched on the consequences of tool selection bias in the introduction. In this section, we will further elaborate our points on why tool selection bias matters to give the reader a comprehensive view of the the effects and significance of the bias we uncover.
L.1 Economic Consequences
Problem Statement. A pay-per-request pricing model is common practice within tool marketplaces (see RapidAPI’s pricing model or that of BridgeAPI). This means that usage of an API directly corresponds to the amount of revenue the developer of that API makes. Ideally, when two or more tools offer identical functionality, usage (and thus revenue) should also be split equally. This is not achieved when LLMs show consistent bias among functionality equivalent tools. This is not a hypothetical concern. As we show in our experiments, some tools gets selected up 10 times more than others whilst being functionally identical (see the geocoding cluster in Figure 16). If revenue is directly correlated with usage, the developer of this tool can then also expect a 10 times higher revenue than the developer of the disadvantaged tool for seemingly no particular reason but more advantageous metadata phrasing (see section 4.3 investigating why tool-selection bias occurs).
Magnitude. The size of the tool-usage market is hard to determine exactly. RapidAPI alone handles over 9 billion requests per month and recommends \$0.00003 per API call at a minimum, a quick calculation gives us at least \$270.000 in developer revenue per month or \$3.24 million per year on this platform alone (RapidAPI (2025a)). Looking more broadly, the global API marketplace market size was estimated at \$18.00 billion in 2024 and is projected to reach \$49.45 billion by 2030, growing at a CAGR of 18.9% from 2025 to 2030 (GrandViewResearch (2025)). From this, one can see that even small shifts in automated traffic can have a meaningful economic impact. It’s unclear how much of this API traffic is currently due to LLM agents, but with the increasing ubiquity of agents, we expect this share to rise significantly. This makes tool selection bias a tangible and pressing economic issue.
L.2 User Experience
Tool-selection bias can directly degrade user experience when an LLM consistently favors an API that is objectively slower, less accurate, or more costly than its functionally equivalent alternatives. In such scenarios, end users may experience higher latency, lower-quality responses, or unnecessary costs, despite the existence of equally capable tools that would have delivered better performance. A more balanced usage distribution across equivalent APIs mitigates these issues by ensuring that no single suboptimal tool is disproportionately selected purely due to incidental metadata or positional biases. This leads to more stable, predictable, and higher-quality user outcomes.
L.3 Safety & Reliability
Biased tool selection magnifies the system’s vulnerability to manipulated metadata and adversarial tools. Recent work on metadata-poisoning attacks (Mo et al. (2025)) shows that adversaries can craft strategically misleading tool names or descriptions to lure LLM agents into invoking harmful or unreliable APIs. When an LLM already exhibits strong, unintended preferences toward superficial metadata cues, such attacks become significantly easier to execute. In this sense, selection bias is not merely an efficiency problem. It increases the surface area for adversarial exploitation and directly undermines the reliability of agentic systems. A more uniform or semantics-invariant selection process would provide a stronger baseline defense by reducing the influence of manipulable metadata.
L.4 Erosion of Trust & Ecosystem Effects
A further systemic consequence of tool-selection bias is the gradual erosion of trust in API marketplaces. If developers observe that LLM-mediated traffic does not meaningfully reflect the functional quality of their tools but instead hinges on arbitrary metadata preferences, they may perceive the marketplace as unfair or unpredictable. This creates incentives to bypass marketplaces entirely by hard-coding specific APIs into their applications. This outcome undermines the value proposition of marketplaces as neutral, competitive intermediaries. In the long term, this can reduce innovation, further distort competition, and create fragile ecosystems where a small number of arbitrarily preferred tools dominate traffic. Addressing tool-selection bias is therefore critical not only for individual user or developer outcomes, but for maintaining trust, participation, and healthy competition in the broader API economy.